# Rethinking industrial artificial intelligence: a unified foundation framework

### v1

@@ Abstract
Recent advancement in industrial artificial intelligence (AI) is reshaping the industry, driving smarter manufacturing, predictive maintenance, and intelligent decision-making. However, existing approaches often focus primarily on algorithms and models, overlooking the importance of systematically integrating domain knowledge, data, and models to ensure more comprehensive and effective AI solutions. Therefore, the effective development and deployment of Industrial AI solutions require a more comprehensive and systematic approach. To address this gap, this paper summarizes previous research and rethinks the role of industrial AI and presents a unified industrial AI foundation framework comprising three core modules: knowledge module, data module, and model module. These modules help to extend and enhance the industrial AI methodology platform, supporting various industrial applications. In addition, a case study on rotating machinery diagnosis demonstrates the framework’s effectiveness, and several future directions are highlighted for the development of the industrial AI foundation framework.
@@ 1 Introduction
The rapid advancement of industrial artificial intelligence (AI) is reshaping industries worldwide [24, 29, 32, 43]. Recent breakthroughs in technologies such as deep learning [24, 29], industrial internet of things (IIoT) [48, 25], large language models (LLMs) [31, 9], prognostics and health management [71, 52], big data analytics [63, 59], and cyber-physical systems (CPS) [44, 28] have accelerated the adoption of industrial AI, enabling industrial systems to extract actionable insights from vast amounts of industrial data and support intelligent decision-making. However, current approaches often overemphasize algorithms and models and often lack a unified framework that systematically integrates domain knowledge, data, and models. Therefore, to unlock the full potential of industrial AI, a structured framework is needed – one that integrates domain knowledge, high-quality data, and intelligent AI models to address complex problems in real-world industrial settings.
Recognizing this gap, this paper presents a unified industrial AI foundation framework composed of knowledge, data, and model modules, enhancing the industrial AI methodology platform. The remainder of the paper is structured as follows: Section 2 reviews related work; Section 3 redefines the role of industrial AI; Section 4 introduces the proposed foundation framework; Section 5 provides a case study on rotating machinery diagnosis; Section 6 discusses future directions; and Section 7 concludes the paper.
Table 1: Summary of existing frameworks for industrial AI applications.
Authors Proposed industrial AI frameworks
Lee et al.
(2018) [29]
Proposes an industrial AI ecosystem framework integrating enabling technologies—data technology (DT), analytic technology (AT),
platform technology (PT), and operation technology (OT)—within the Cyber-Physical System (5C) architecture. The framework
systematically guides AI implementation in smart manufacturing under the guidance of five key elements: analytics technology, big
data technology, cloud or cyber technology, domain knowhow, and evidence (ABCDE).
Zhang et al.
(2019) [68]
Presents a comprehensive Industrial AI reference framework consisting of seven key dimensions: object (who), domain (where),
@@
Peres et al.
(2020) [43]
Proposes a conceptual Industrial AI framework highlighting essential enabling technologies (data, analytics, platforms, operations,
and human-machine interaction), while systematically identifying critical challenges attributes, capabilities, design principles, and
@@
Yang et al.
(2021) [65]
Presents a two-tier intelligent manufacturing framework designed for the process industry, integrating human-machine cooperation
and intelligent autonomous control systems, to achieve intelligent optimal decision-making.
Ahmed et al.
(2022) [2]
Conducts a comprehensive survey on AI and explainable artificial intelligence (XAI) methodologies within Industry 4.0,
categorizing various XAI approaches (model-specific, model-agnostic, local/global, visualization-based methods, and surrogate
models), highlighting their applicability, benefits, and challenges in industrial contexts.
Jan et al.
(2023) [24]
Proposes a structured, four-stage industrial AI data pipeline (data acquisition/validation, data processing/fusion, model training/testing,
and model interpretation). Identifies common themes, issues, and industry-specific solutions related to AI integration in various sectors,
providing insights into practical challenges and opportunities within Industry 4.0.
Leng et al.
(2024) [32]
Presents a technical reference framework structured in four layers (hardware infrastructure, computing engine, AI algorithms, and
industrial application layer together with related empowering technologies across layers), identifying three core opportunities
(collaborative intelligence, self-learning intelligence, crowd intelligence) crucial for realizing Industry 5.0’s vision of human-centric,
resilient, and sustainable manufacturing.
Lee and Su
(2024) [31]
Proposes a unified industrial large knowledge model (ILKM) framework consisting of four systematic steps: (1) construction of a large
knowledge library, (2) preparation of domain-specific instruction data, (3) development of domain-specific knowledge large language
models, and (4) establishment of intelligent domain expert machine learning systems, guided by the ”6S Principle” (specialized
knowledge, scrutability, safety, scalability, sustainability, systematization).
@@ 2 Related work
This section reviews existing studies and frameworks focusing on the integration of AI within various industrial applications. Several research works have proposed frameworks aiming to enhance industrial productivity, decision-making, and operations by leveraging AI capabilities. Table 1 provides a clear summary of these previous industrial AI frameworks, highlighting their primary contributions. For instance, Lee et al. introduced an industrial AI ecosystem framework integrating enabling technologies such as data technology, analytic technology, platform technology, and operation technology within the established CPS (5C) architecture, guided by key ABCDE elements [29]. Building upon this foundation, Peres et al. expanded the framework into a more comprehensive conceptual framework, emphasizing challenges, design principles, essential technologies, capabilities, attributes, and industrial application areas [43]. Zhang et al. developed a comprehensive reference framework structured around seven critical dimensions (object, domain, stage, requirement, technology, function, and solutions) [68]. In addition, Yang et al. proposed an intelligent manufacturing framework combining human-machine cooperation with autonomous intelligent control systems specifically tailored to the process industry [65]. Ahmed et al. highlighted AI and explainable artificial intelligence (XAI) methodologies, categorizing various XAI methods and their applicability in Industry 4.0 contexts [2], while Jan et al. structured a four-stage AI data pipeline, outlining data acquisition/validation, data processing/fusion, model training/testing, and model interpretation [24]. Furthermore, Leng et al. introduced a four-layer technical reference framework encompassing layers from hardware to industrial applications [32]. Most recently, Lee and Su proposed a unified industrial large knowledge model (ILKM) framework emphasizing domain-specific knowledge integration through LLMs and machine learning approaches [31].
Although these existing frameworks have successfully identified important aspects of AI integration across different industrial applications, several limitations remain. These include the lack of systematic domain-specific knowledge integration, a tendency to focus on theoretical conceptualization without clear and objective guidelines, and an overemphasis on data and models without sufficient attention to systematic thinking that connects knowledge, data, and models. Motivated by these gaps, the following two sections revisit the role of industrial AI and present a unified industrial AI foundation framework aimed at addressing these critical shortcomings.
@@ 3 Rethinking the role of industrial AI
The role of industrial AI can be best understood through a two-layer perspective, as shown in Figure 1, where the physical layer consists of human, things, and systems, while the digital layer consists of knowledge, data, and models. Based on recent emerging technologies such as IIoT [48, 25], digital twin [54, 16], CPS [44, 28], industrial big data analytics [63, 59], deep learning [24, 29], and LLMs [31, 9], industrial AI acts as a dynamic bridge between the physical and digital layers, continuously refining and applying AI-driven insights to improve and optimize the real-world industrial systems. More specifically, human expertise forms the basis of domain knowledge, which contains accumulated experience, research insights, and best practices from AI-driven approaches. With the rise of IoT technology, things, including machines, sensors, and other devices, generate data, and data is the essential prerequisite to conducting industrial AI analytics, modeling, and decision-making in modern industry. Moreover, systems such as CPS, IIoT systems, and other intelligent industrial systems use AI models to automate operations, enable intelligent decision-making, and improve real-time control.
While the two-layer perspective provides a conceptual understanding of how industrial AI bridges the physical and digital domains, real-world industrial settings present several challenges. In practice, developers and engineers face persistent difficulties in transforming heterogeneous data from IIoT devices into reliable, AI-ready formats due to inconsistent data quality, lack of standard preprocessing pipelines, and fragmented metadata management [48]. CPS, although widely discussed, often remains difficult to operationalize, as maintaining consistency between real-world operations and virtual models requires continuous data synchronization and robust model updating [44]. Digital twin development, despite its promise, is challenged by the complexity of integrating multi-source data streams, real-time analytics, and simulation modeling requirements [54]. Furthermore, the increasing availability of LLMs introduces new opportunities but also poses practical challenges in domain knowledge understanding, extraction, and interpretability for industrial use cases [31]. Based on our observations from industry collaborations and project experiences, these complexities frequently result in ad-hoc solutions, isolated development efforts, and inefficiencies in scaling AI-driven applications. Therefore, these observations and challenges motivate the need for a structured foundation framework that systematically connects knowledge, data, and models, as presented in the following section.
Figure 1: The role of industrial AI: bridging the physical and digital layers.
@@ 4 Industrial AI foundation framework
In this section, we propose a unified Industrial AI foundation framework designed to systematically guide the development and deployment of industrial AI solutions, as illustrated in Figure 2. The framework consists of three core modules: (i) Knowledge module; (ii) Data module; and (iii) Model module. Within each module, we identify four key components, resulting in twelve important aspects that collectively provide structured direction for industrial AI developers and practitioners. These modules enhance and extend the Industrial AI methodology platform by providing a structured, modular foundation for existing and emerging industrial AI methodologies. Figure 2 illustrates the overall structure of the proposed framework, highlighting the interconnections and dynamic feedback between knowledge, data, and model modules. Each pair of modules is connected by forward and backward arrows, highlighting that these modules continuously inform and refine one another, rather than existing in a linear sequence. In addition, forward arrows from all three modules point toward the industrial AI methodology platform, indicating that the platform is strengthened by these three modules. In the subsequent Sections 4.1–4.4, the details of the knowledge, data, and model modules are outlined, followed by an explanation of how they enhance and extend the capabilities of the industrial AI methodology platform.
@@ 4.1 Knowledge module
Knowledge module focuses on capturing, organizing, and utilizing domain knowledge to support the entire industrial AI development cycle. In complex industrial settings, domain expertise often exists in unstructured forms such as reports, manuals, research papers, and maintenance logs. Rather than reflecting an anthropocentric design philosophy in the traditional sense, the knowledge module aims to transform these scattered resources into structured, accessible, and reusable formats that guide data analytics and model development—serving as a foundation that enables scalable, automated, and context-aware AI system development. It focuses on the following four aspects.
@@ 4.1.4 New knowledge generation and integration
Figure 2: A unified industrial AI foundation framework.
@@ 4.2.1 Data preprocessing
This is a fundamental prerequisite for successful Industrial AI applications, as raw data often contains issues such as noise, missing values, anomalies, class imbalances, and labeling inconsistencies. A well-structured data preprocessing pipeline typically involves several key techniques, including outlier detection methods (such as isolation forests, local outlier factor, and statistical thresholding); data imputation methods (including mean, median, multiple imputation, K-nearest neighbor, and regression-based approaches); signal denoising methods (such as wavelet transforms and filtering); and resampling methods (including synthetic minority over-sampling technique, random over-sampling, and under-sampling) [11]. In addition, researchers are encouraged to adopt, design and document reusable preprocessing pipelines using standardized tools such as Pandas, scikit-learn, PyTorch, and TensorFlow to ensure reproducibility and scalability.
@@ 4.2.2 Data quality
Data quality directly impacts the reliability of AI models, as poor-quality data can degrade model performance and lead to poor decision-making. Key dimensions of data quality include consistent representation, completeness, uniqueness, feature accuracy, target accuracy, and target class balance [8]. Therefore, developers should adopt systematic data quality checks as part of their pipelines. For instance, “CleanML” demonstrates how various data quality issues can significantly affect the performance of common machine learning models [34]. Furthermore, Foroni et al. extend conventional data quality definitions by evaluating not only how data deviates from an ideal clean dataset but also how these deviations influence task outcomes [17].
@@ 4.2.3 Feature engineering
Feature engineering focuses on extracting meaningful representations from raw data, improving model performance, interpretability, and efficiency [69]. In industrial applications, well-designed features can greatly improve model accuracy. Common techniques include domain-specific feature extraction, such as time-domain statistical measures, frequency-domain features derived from Fourier transforms, and time-frequency domain features. Dimensionality reduction methods, including principal component analysis (PCA), t-distributed stochastic neighbor embedding (t-SNE), and linear discriminant analysis, help reduce complexity while maintaining important information. Feature importance ranking methods are used to identify key variables that most influence model predictions, guiding feature selection. In addition, deep learning-based feature representation learning can discover complex patterns in datasets. Tools such as tsfresh, PyCaret, and SHAP provide automated pipelines for feature extraction, importance, and selection. Therefore, effective feature engineering ensures that AI models focus on the most informative inputs while minimizing redundancy and noise.

### v2

@@ Abstract
Recent advancements in industrial artificial intelligence (AI) are reshaping the industry by driving smarter manufacturing, predictive maintenance, and intelligent decision-making. However, existing approaches often focus primarily on algorithms and models while overlooking the importance of systematically integrating domain knowledge, data, and models to develop more comprehensive and effective AI solutions. Therefore, the effective development and deployment of industrial AI require a more comprehensive and systematic approach. To address this gap, this paper reviews previous research, rethinks the role of industrial AI, and proposes a unified industrial AI foundation framework comprising three core modules: the knowledge module, data module, and model module. These modules help to extend and enhance the industrial AI methodology platform, supporting various industrial applications. In addition, a case study on rotating machinery diagnosis is presented to demonstrate the effectiveness of the proposed framework, and several future directions are highlighted for the development of the industrial AI foundation framework.
@@ 1 Introduction
The rapid advancement of industrial artificial intelligence (AI) is reshaping industries worldwide. [23, 29, 32, 43]. Recent breakthroughs in technologies such as deep learning [23, 29], industrial internet of things (IIoT) [48, 24], large language models (LLMs) [31, 9], prognostics and health management [71, 52], big data analytics [63, 59], and cyber-physical systems (CPS) [44, 28] have accelerated the adoption of industrial AI, enabling industrial systems to extract actionable insights from vast amounts of industrial data and support intelligent decision-making. However, current approaches often overemphasize algorithms and models while lacking a unified framework that systematically integrates domain knowledge, data, and models. To unlock the full potential of industrial AI, a structured framework is needed – one that integrates domain knowledge, high-quality data, and intelligent AI models to address complex challenges in real-world industrial settings
Recognizing this gap, this paper proposes a unified industrial AI foundation framework composed of knowledge, data, and model modules to enhance the industrial AI methodology platform. The remainder of the paper is structured as follows: Section 2 reviews related work; Section 3 redefines the role of industrial AI; Section 4 introduces the proposed foundation framework; Section 5 provides a case study on rotating machinery diagnosis; Section 6 discusses future directions; and Section 7 concludes the paper.
Table 1: Summary of existing frameworks for industrial artificial intelligence applications
Authors Year Proposed industrial artificial intelligence (AI) frameworks
Lee et al. [29]
2018
Proposed an industrial AI ecosystem framework to integrate enabling technologies – data technology, analytic technology, platform
technology, and operation technology – within the Cyber-Physical System (5C) architecture. The framework systematically guides
AI implementation in smart manufacturing under the guidance of five key ABCDE elements: analytics technology, big data
technology, cloud or cyber technology, domain know-how, and evidence.
Zhang et al. [68]
2019
Presented a comprehensive industrial AI reference framework consisting of seven key dimensions: Object (who), domain (where),
@@
Peres et al. [43]
2020
Proposed a conceptual industrial AI framework highlighting essential enabling technologies (data, analytics, platforms, operations,
and human-machine interaction) while systematically identifying critical challenges, attributes, capabilities, design principles, and
@@
Yang et al. [65]
2021
Presented a two-tier intelligent manufacturing framework designed for the process industry, integrating human-machine cooperation,
and intelligent autonomous control systems to achieve intelligent optimal decision-making.
Ahmed et al. [2]
2022
Conducted a comprehensive survey on AI and explainable AI (XAI) methodologies within Industry 4.0, categorizing various XAI
approaches (model-specific, model-agnostic, local/global, visualization-based methods, and surrogate models), highlighting their
applicability, benefits, and challenges in industrial contexts.
Jan et al. [23]
2023
Proposed a structured, four-stage industrial AI data pipeline (data acquisition/validation, data processing/fusion, model training/testing,
and model interpretation). They identified common themes, issues, and industry-specific solutions related to AI integration in
various sectors, providing insights into practical challenges and opportunities within Industry 4.0.
Leng et al. [32]
2024
Presented a technical reference framework structured in four layers (hardware infrastructure, computing engine, AI algorithms,
and industrial application layer together with related empowering technologies across layers), identifying three core opportunities
(collaborative intelligence, self-learning intelligence, and crowd intelligence) crucial for realizing Industry 5.0’s vision of
human-centric, resilient, and sustainable manufacturing.
Lee and Su [31]
2024
Proposed a unified industrial large knowledge model framework consisting of four systematic steps: (i) construction of a large
knowledge library; (ii) preparation of domain-specific instruction data; (iii) development of domain-specific knowledge large language
models; and (iv) establishment of intelligent domain expert machine learning systems; guided by the “6S Principle” (specialized
knowledge, scrutability, safety, scalability, sustainability, and systematization).
@@ 2 Related work
This section reviews existing studies and frameworks focusing on the integration of AI within various industrial applications. Several research works have proposed frameworks aimed at enhancing industrial productivity, decision-making, and operations by leveraging AI capabilities. Table 1 provides a clear summary of these previous industrial AI frameworks, highlighting their primary contributions. For instance, Lee et al. [29] introduced an industrial AI ecosystem framework that integrates enabling technologies, such as data technology, analytic technology, platform technology, and operation technology, within the established CPS (5C) architecture, guided by key ABCDE elements. Building on this foundation, Peres et al. [43] expanded the framework into a more comprehensive conceptual framework, emphasizing challenges, design principles, essential technologies, capabilities, attributes, and industrial application areas. On the other hand, Zhang et al. [68] developed a comprehensive reference framework structured around seven critical dimensions (object, domain, stage, requirement, technology, function, and solutions). In addition, Yang et al. [65] proposed an intelligent manufacturing framework combining human-machine cooperation with autonomous intelligent control systems specifically tailored to the process industry. Ahmed et al. [2] highlighted AI and explainable AI (XAI) methodologies, categorizing various XAI methods and their applicability in Industry 4.0 contexts, while Jan et al. [23] structured a fourstage AI data pipeline, outlining data acquisition/validation, data processing/fusion, model training/testing, and model interpretation. Furthermore, Leng et al. [32] introduced a fourlayer technical reference framework encompassing layers from hardware to industrial applications. Most recently, Lee and Su [31] proposed a unified industrial large knowledge model (ILKM) framework emphasizing domain-specific knowledge integration through LLMs and machine learning (ML) approaches.
Although these frameworks have successfully identified important aspects of AI integration across different industrial applications, several limitations remain. These include the lack of systematic domain-specific knowledge integration, a tendency to focus on theoretical conceptualization without clear and objective guidelines, and an overemphasis on data and models without sufficient attention to systematic thinking that connects knowledge, data, and models. Motivated by these gaps, the following two sections revisit the role of industrial AI and present a unified industrial AI foundation framework aimed at addressing these critical shortcomings.
@@ 3 Rethinking the role of industrial AI
The role of industrial AI can be best understood through a two-layer perspective, as shown in Figure 1, where the physical layer consists of humans, things, and systems, while the digital layer consists of knowledge, data, and models. Based on recent emerging technologies such as IIoT [48, 24], digital twin [54, 27], CPS [44, 28], industrial big data analytics [63, 59], deep learning [23, 29], and LLMs [31, 9], industrial AI acts as a dynamic bridge between the physical and digital layers, continuously refining and applying AI-driven insights to improve and optimize the real-world industrial systems. More specifically, human expertise forms the basis of domain knowledge, which contains accumulated experience, research insights, and best practices from AI-driven approaches. With the rise of IoT technology, things, including machines, sensors, and other devices, generate vast amounts of data, which is an essential prerequisite for conducting industrial AI analytics, modeling, and decision-making in modern industry. Moreover, systems such as CPS, IIoT systems, and other intelligent industrial systems use AI models to automate operations, enable intelligent decision-making, and improve real-time control.
While the two-layer perspective provides a conceptual understanding of how industrial AI bridges the physical and digital domains, real-world industrial settings present several challenges. In practice, developers and engineers face persistent difficulties in transforming heterogeneous data from IIoT devices into reliable, AI-ready formats due to inconsistent data quality, lack of standard preprocessing pipelines, and fragmented metadata management [48]. CPS, although widely discussed, often remains difficult to operationalize, as maintaining consistency between realworld operations and virtual models requires continuous data synchronization and robust model updating [44]. Despite its promise, digital twin development is challenged by the complexity of integrating multi-source data streams, real-time analytics, and simulation modeling requirements [54]. Furthermore, the increasing availability of LLMs introduces new opportunities but also pose practical challenges in domain knowledge understanding, extraction, and interpretability for industrial use cases [31]. Based on our observations from industry collaborations and project experiences, these complexities frequently result in ad-hoc solutions, isolated development efforts, and inefficiencies in scaling AI-driven applications. Therefore, these observations and challenges motivate the need for a structured foundation framework that systematically connects knowledge, data, and models, as presented in the following section.
Figure 1: The role of industrial artificial intelligence in bridging the physical and digital layers. Image created by the authors.
@@ 4 Industrial AI foundation framework
In this section, we propose a unified industrial AI foundation framework designed to systematically guide the development and deployment of industrial AI solutions, as illustrated in Figure 2. The framework consists of three core modules: (i) knowledge module; (ii) data module; and (iii) model module. Within each module, we identify four key components, resulting in 12 important aspects that collectively provide structured direction for industrial AI developers and practitioners. These modules enhance and extend the industrial AI methodology platform by providing a structured, modular foundation for existing and emerging industrial AI methodologies. Figure 2 illustrates the overall structure of the proposed framework, highlighting the interconnections and dynamic feedback between knowledge, data, and model modules. Each pair of modules is connected by forward and backward arrows, highlighting that these modules continuously inform and refine one another, rather than existing in a linear sequence. In addition, forward arrows from all three modules point toward the industrial AI methodology platform, indicating that the platform is strengthened by these three modules. In Sections 4.1-4.4, the details of the knowledge, data, and model modules are outlined, followed by an explanation of how they enhance and extend the capabilities of the industrial AI methodology platform.
@@ 4.1 Knowledge module
The knowledge module focuses on capturing, organizing, and utilizing domain knowledge to support the entire industrial AI development cycle. In complex industrial settings, domain expertise often exists in unstructured forms, such as reports, manuals, research papers, and maintenance logs. Rather than reflecting an anthropocentric design philosophy in the traditional sense, the knowledge module aims to transform these scattered resources into structured, accessible, and reusable formats that guide data analytics and model development – serving as a foundation that enables scalable, automated, and context-aware AI system development. It focuses on the following four aspects.
@@ 4.1.4 New knowledge generation and integration
Figure 2: A unified industrial artificial intelligence (AI) foundation framework. The industrial large knowledge model framework image is reprinted with permission from Lee and Su [31]. The stream of quality image is reprinted with permission from Lee et al. [30] Copyright © 2022 Society of Manufacturing Engineers (SME). The 5C-cyber-physical system figure is reprinted with permission from Lee et al. [29] Copyright © 2018 SME. The digital twin image is reprinted with permission from Lee et al. [27] Copyright © 2020 The Institution of Engineering and Technology. Abbreviations: LLM: Large language model; ML: Machine learning.
@@ 4.2.1 Data preprocessing
This is a fundamental prerequisite for successful industrial AI applications, as raw data often contains issues such as noise, missing values, anomalies, class imbalances, and labeling inconsistencies. A well-structured data preprocessing pipeline typically involves several key techniques, including outlier detection methods (such as isolation forests, local outlier factor, and statistical thresholding); data imputation methods (including mean, median, multiple imputation, K-nearest neighbors, and regression-based approaches); signal denoising methods (such as wavelet transforms and filtering); and resampling methods (including synthetic minority over-sampling technique, random over-sampling, and under-sampling) [11]. In addition, researchers are encouraged to adopt, design and document reusable preprocessing pipelines using standardized tools such as Pandas, scikit-learn, PyTorch, and TensorFlow to ensure reproducibility and scalability.
@@ 4.2.2 Data quality
Data quality directly impacts the reliability of AI models, as poor-quality data can degrade model performance and lead to poor decision-making. Key dimensions of data quality include consistent representation, completeness, uniqueness, feature accuracy, target accuracy, and target class balance [8]. Therefore, developers should adopt systematic data quality checks as part of their pipelines. For instance, CleanML demonstrates how various data quality issues can significantly affect the performance of common machine learning models [34]. Furthermore, Foroni et al. [16] extend conventional data quality definitions by evaluating not only how data deviates from an ideal clean dataset but also how these deviations influence task outcomes.
@@ 4.2.3 Feature engineering
Feature engineering focuses on extracting meaningful representations from raw data, improving model performance, interpretability, and efficiency [69]. In industrial applications, well-designed features can greatly improve model accuracy. Common techniques include domain-specific feature extraction, such as time-domain statistical measures, frequency-domain features derived from Fourier transforms, and time-frequency domain features. Dimensionality reduction methods, including principal component analysis (PCA), t-distributed stochastic neighbor embedding (t-SNE), and linear discriminant analysis, help reduce complexity while maintaining important information. Feature importance ranking methods are used to identify key variables that most influence model predictions, guiding feature selection. In addition, deep learning-based feature representation learning can discover complex patterns in datasets. Tools such as tsfresh, PyCaret, and Shapley Additive Explanations (SHAP) provide automated pipelines for feature extraction, importance, and selection. Therefore, effective feature engineering ensures that AI models focus on the most informative inputs while minimizing redundancy and noise.